please help me solve this problem , thank you .
Ten years ago, a father was three times as old as his sons. If he is now only twice as old as his son, find their present ages.

Answer by stanbon(75887)

Ten years ago, a father was three times as old as his son. If he is now only twice as old as his son, find their present ages.
----
Equations:
(f-10) = 3(s-10)
f = 2s
-----
Substitute for "f" and solve for "s":
(2s-10) = 3s-30
s = 20 yrs (Son's age now)
----
Solve for "f":
f = 2s
f = 40 yrs (father's age now)
=================================
Cheers,
Stan H.
===========